### v1

# QoS-aware State-Augmented Learnable Algorithm for Wireless Coexistence Parameter Management

Abstract
Efficient and fair coexistence in unlicensed spectrum is essential to support heterogeneous networks such as 5G NR-U and Wi-Fi, which often contend for shared wireless resources. We introduce a general framework for wireless Coexistence Parameter Management (CPM) based on state-augmented constrained reinforcement learning. We propose a novel algorithm, QaSAL-CPM, which incorporates state augmentation by embedding the dual variables in the constrained optimization formulation directly into the agent’s observation space. This method enables the agent to respond to constraint violations in real time while continuing to optimize a primary performance objective. Through extensive simulations of 5G NR-U and Wi-Fi coexistence scenarios, we show that QaSAL-CPM achieves reliable QoS compliance and improved policy robustness across various transmitter densities compared to previous approaches. The proposed framework offers a scalable and adaptive solution for real-time coexistence optimization in next-generation wireless networks.
Index Terms:
5G NR-U, Wi-Fi, coexistence, unlicensed spectrum, quality of service (QoS), medium access delay, fairness.
I Introduction
[…]
[findent=2pt]The rapid growth in wireless connectivity demands driven by applications ranging from mobile broadband to the Internet of Things (IoT) has led to increased reliance on unlicensed spectrum. These frequency bands are shared by multiple Radio Access Technologies (RATs), such as 5G New Radio in Unlicensed Spectrum (NR-U) and Wi-Fi, creating complex interference dynamics that, if left unmanaged, can significantly degrade network performance. The key challenge lies in ensuring high performance while maintaining fairness among the coexisting technologies sharing the same spectrum [1, 2, 3].
The coexistence of 5G NR-U and Wi-Fi introduces unique difficulties due to their distinct medium access protocols. While NR-U relies on Listen-Before-Talk (LBT), Wi-Fi uses Carrier Sense Multiple Access with Collision Avoidance (CSMA/CA). These different protocols should coordinate effectively to minimize inter-network collisions and achieve fair spectrum utilization. The situation is further complicated by divergent Quality of Service (QoS) requirements and traffic behaviors. For instance, 5G NR-U often supports latency-sensitive, high-priority services that may conflict with Wi-Fi’s best-effort traffic model. Moreover, the dynamic nature of unlicensed spectrum usage—due to fluctuating traffic loads, mobility, and environmental changes—demands adaptive and robust coexistence strategies [4]. Without such strategies, the performance of both technologies can deteriorate due to increased contention, latency, and unfair access. Managing contention and enhancing performance under these conditions can be formulated as a QoS-aware network utility maximization problem, in which the tuning of Medium Access Control (MAC) parameters (e.g., contention window sizes), which we refer to as Coexistence Parameter Management (CPM), is key.
Multi-objective reinforcement learning (RL) has been applied to this problem by defining rewards as weighted combinations of performance metrics such as delay and fairness [5, 6]. However, this approach requires manually selecting and tuning weights, which introduces substantial computational overhead and does not guarantee strict adherence to QoS requirements. In contrast, constrained RL explicitly incorporates QoS constraints into the learning process [7]. Constrained RL formulates the problem using Lagrangian relaxation, where a single objective function is optimized while constraints are enforced through dynamic dual variables. This method automatically adjusts the Lagrange multipliers, simplifying the design and enabling better constraint satisfaction over time. Still, dual-based methods can struggle in highly dynamic environments due to difficulties in tuning penalty terms and ensuring stable convergence [8, 9].
In this paper, we propose a general framework for CPM based on a state-augmented RL formulation, which was developed recently in [10, 11] to overcome the drawbacks of the dual-based methods. In state-augmented RL, the dual variables are embedded directly into the agent’s observation space, enabling real-time responsiveness to constraint violations and reducing reliance on hand-tuned penalties. Building upon our earlier works [12, 13], which introduced traffic-aware RL and the application of state-augmented RL to the wireless coexistence problem, we present a new algorithm, QaSAL-CPM, which integrates state augmentation with a Double Deep Q-Network (DDQN) and experience replay. QaSAL-CPM achieves real-time QoS guarantees for high-priority traffic while maintaining fairness across all traffic classes in a wireless coexistence scenario. We implemented a Python-based simulation environment that models the MAC-layer behavior of 5G NR-U and Wi-Fi under saturated traffic conditions to compare the performance of QaSAL-CPM to that of multi-objective RL and constrained RL based on conventional primal-dual methods. The proposed QaSAL-CPM provides a framework for significantly improved QoS compliance and policy robustness across diverse wireless coexistence scenarios.
The rest of the paper is organized as follows: Section II reviews related work on general coexistence strategies, the application of RL to NR-U/Wi-Fi coexistence, and constrained RL techniques. Section III presents the system model and formally defines the CPM problem. In Section IV, we develop a multi-objective RL approach for CPM, which extends our prior work in [12]. Section V introduces a constrained RL formulation, discusses a primal-dual method for solving it, and presents the state-augmented RL framework. This motivates the design of our proposed algorithm, QaSAL-CPM, which is detailed in Section VI along with its deep RL implementation. Section VII applies the CPM framework to the coexistence of 5G NR-U and Wi-Fi in unlicensed spectrum, discusses the performance metrics of interest, the simulation setup, and the numerical results. The paper is concluded in Section VIII.
II Related Works
II-A General Coexistence Studies
Coexistence of 5G NR-U/LAA and WiFi has been studied extensively in the recent literature. For example, [1] and [2] provide comprehensive overviews of the state-of-the-art research on coexistence of different technologies in sub-7 GHz unlicensed spectrum. A detailed overview of the new 5G NR-U specifications and its coexistence with Wi-Fi in unlicensed bands, focusing on key challenges in PHY, MAC, and higher-layer operations is presented in [3]. A comprehensive survey on the opportunities and challenges associated with unlicensed operations in the 6 GHz band, focusing on Wi-Fi and 5G NR-U coexistence, can be found in [14]. In [15], the coexistence of licensed fixed service systems with unlicensed Radio Local Area Networks and 5G NR-U in the 6 GHz band, is investigated, focusing on potential 5G NR-U and Wi-Fi interference risks.
II-B Bandit-Based Reinforcement Learning
The Multi-Armed Bandit (MAB) problem is a fundamental reinforcement learning framework that models decision-making under uncertainty, where an agent must choose among multiple arms to maximize cumulative rewards over time. In each round, the agent selects an action, observes a reward, and updates its strategy to balance exploration (gathering new information) and exploitation (leveraging past knowledge). In the context of coexistence management for unlicensed spectrum, MAB algorithms provide an adaptive mechanism for optimizing channel access, power control, and contention window adjustments in dynamic environments where 5G NR-U and Wi-Fi compete for shared resources.
In [16], an MAB-based online learning distributed channel selection algorithm is proposed to enable NR-U users to make optimal channel selections without requiring complete environmental knowledge. An AI-driven framework is introduced in [17] for adaptive sensing threshold selection in shared spectrum environments by leveraging a clustering-based multi-armed bandit algorithm, ensuring efficient coexistence among Wi-Fi, LTE-LAA, and 5G NR-U systems. A novel framework for fair and efficient NR-U coexistence in shared spectrum environments is presented in [18], which formulates the channel access problem as a multi-objective MAB that jointly optimizes spectrum efficiency and fairness.
II-C State-Driven Reinforcement Learning
MAB algorithms operate under a context-free learning paradigm, where decisions are made without considering past states or future consequences. In real-world coexistence scenarios, however, the wireless environment is highly dynamic—interference patterns, traffic loads, and network conditions fluctuate over time, requiring an approach that can learn and adapt based on these changes. State-driven RL methods address this limitation of MAB-based approaches by incorporating state transitions and delayed rewards, which facilitates long-term optimization of coexistence strategies.
Figure 1: A 5G NR-U/Wi-Fi coexistence in an unlicensed spectrum with multi-priority traffic transmitter.
In [19], an RL-based approach is proposed for joint allocation of transmission opportunities and spectral resources in 5G NR-U and Wi-Fi coexistence while ensuring fair spectrum sharing. In [20], a joint base station and resource allocation framework based on deep RL is presented for a heterogeneous network with vehicle-to-everything users. Deep RL-based approaches for energy detection threshold optimization in NR-U and Wi-Fi coexistence systems, focusing on downlink Ultra-Reliable Low-Latency Communications (URLLC), are proposed in [21]. In [22], the challenge of fair coexistence between 5G NR and Wi-Fi while accommodating URLLC traffic is addressed through a mixed priority scheduling mechanism that aims to balance latency constraints and data size. The previous RL-based approaches do not explicitly address QoS constraints within the wireless coexistence context, which is the main contribution of our work.
III System Model and Problem Formulation
III-A System Model
We consider a 5G NR-U network coexisting with Wi-Fi networks on an unlicensed frequency band as illustrated in Fig. 1. Transmitters of both networks contend with each other to access the channel for downlink transmissions. We focus on a worst-case saturated scenario, where transmitters in both networks always have packets to transmit. To access the channel, the Wi-Fi network adopts the IEEE 802.11 enhanced distributed channel access (EDCA) protocol, which uses binary exponential backoff with a random initial backoff value. EDCA defines four access classes, each of which has a different initial backoff window size, maximum backoff stages, and maximum transmission opportunity (TXOP) duration.
NR-U gNB transmitters, on the other hand, employ the LBT mechanism, which is consistent with Wi-Fi’s EDCA, and supports four access priority classes that correspond to the four access classes in EDCA [23]. For the PHY layer, gNBs have flexible numerologies and time slot scheduling in which each NR-U transmission can only begin at the next slot boundary. If the channel is found idle by a gNB, it transmits a reservation signal (RS) until the next slot boundary to prevent other transmitters from occupying the channel. For simplicity, we assume that each NR-U transmitter has the same channel occupation times for successful transmissions and collisions as Wi-Fi.
Without loss of generality, we assume two types of transmitters for each network: 1) high-priority traffic transmitters with strict delay requirements and 2) low-priority traffic transmitters with less stringent delay constraints. Throughout this work, we denote the transmitters with high-priority and low-priority traffic by PC1 (priority class 1) and PC3 (priority class 3), respectively. The LBT protocol can harm the latency performance of PC1 traffic such as URLLC packets, especially when contending for channel access with PC3 traffic [24]. Additionally, the LBT protocol in NR-U may cause additional transmission delay compared to 5G NR in licensed spectrum due to the unpredictability of transmission opportunities. Therefore, the delay performance of PC1 traffic may not always meet the requirements.
Effective coexistence management between 5G NR-U and Wi-Fi in unlicensed spectrum requires adaptive parameter tuning to balance fair spectrum access and QoS performance (e.g., latency, efficiency, and throughput) for PC1 traffic transmitters. Examples of key coexistence parameters include the Contention Window (CW) size and the Arbitration Inter-Frame Space (AIFS) duration, which directly impact how transmitters of both networks compete for channel access. CPM refers to the dynamic adjustment of these parameters based on real-time network conditions. The objective of CPM is to ensure the QoS requirements for high-priority traffic while maintaining fair and efficient coexistence. Fig. 2 illustrates the parameterized optimization of the CPM policy through a Double Deep Q-Network (DDQN), which will be discussed further in Section VI.
Figure 2: Illustration of parameterized optimization of the CPM policy through Deep Q-Network (DQN).
III-B Problem Formulation
Let $\mathpzc{S}\subset\mathds{R}^{n}$ represent a compact set of coexistence environment states capturing the status of both the 5G NR-U and Wi-Fi networks. Given the state $\mathbf{S}_{t}\in\mathpzc{S}$, let $\mathbf{a}(\mathbf{S}_{t})$ denote the vector of CPM decisions across both networks, where $\mathbf{a}:\mathpzc{S}\rightarrow\mathds{R}^{a}$ is the CPM function. The possible states and CPM decisions are described by a Markov Decision Process (MDP) with transition probability $p\left(\mathbf{S}_{t+1}|\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t})\right)$. The agent sequentially makes CPM decisions over discrete time steps $t\in\mathds{N}\cup\{0\}$ based on a policy $\pi$, and observes the performance vector $\mathbf{f}(\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t}))\in\mathds{R}^{m}$ which captures various performance metrics of both networks, such as delay, collision rate, airtime efficiency, fairness, number of successful transmissions, channel utilization ratio, etc.
As with any MDP, the primary concern is the network performance the agent accumulates over an infinite horizon, which is represented by the value function
$V_{i}(\pi)\triangleq\lim_{T\rightarrow\infty}\frac{1}{T}\mathds{E}_{\left(\mathbf{S},\mathbf{a}(\mathbf{S})\right)\sim\pi}\left[\sum_{t=0}^{T}f_{i}(\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t}))\right],$ (1)
where $i=0,\ldots,m-1$. In practice, because the probability distribution of the environment is unknown, the objective is derived based on the ergodic average network performance
$\tilde{V}_{i}(\pi)=\frac{1}{T}\sum_{t=0}^{T-1}f_{i}(\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t})).$ (2)
In the coexistence of 5G NR-U and Wi-Fi, multiple conflicting objectives naturally arise due to the need to balance different network performance requirements. Each network operates under distinct access mechanisms and traffic demands, leading to trade-offs between key performance metrics. Some transmitters may require strict QoS guarantees, while others prioritize overall network efficiency or fairness in resource allocation. Optimizing one aspect of the system often comes at the expense of another, making it challenging to achieve an optimal balance through single-objective optimization. This makes CPM a multi-objective optimization problem:
$\max{\tilde{\mathbf{V}}(\pi)}=\max{\left\{\tilde{V}_{0}(\pi),\tilde{V}_{1}(\pi),\dots,\tilde{V}_{m-1}(\pi)\right\}},$ (3)
in which different performance criteria must be optimized together rather than in isolation.
IV Multi-Objective Optimization for CPM
To solve (3), a structured approach is needed to systematically weigh the different objectives and guide the learning process toward solutions that achieve the best possible trade-offs. A common approach to addressing this challenge is scalarization, which applies a function that combines all objectives into a single objective function. This reformulation allows the problem to be treated as a standard single-objective MDP, enabling the agent to learn a policy that maximizes the scalarized objective. A common scalarization approach is linear scalarization, which involves a weight vector $\boldsymbol{w}=\{w_{0},w_{1},\dots,w_{m-1}\}$, $w_{i}\in[0,1]$, $\sum_{i=0}^{m-1}{w_{i}}=1$. Weight $w_{i}$ indicates the relative importance of objective $f_{i}(\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t}))$ in a weighted sum of the objectives:
$f_{\text{scalar}}(\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t}))=\sum_{i=0}^{m-1}{w_{i}f_{i}(\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t}))}.$ (4)
Optimization of $f_{\text{scalar}}(\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t}))$ in (4) involves an infinite-dimensional search over the space of CPM decisions $\mathbf{a}(\mathbf{S})$ for any given network state $\mathbf{S}$, making direct optimization impractical. To address this, a parameterized approach for the CPM policy can be adopted by replacing $\mathbf{a}(\mathbf{S})$ with $\mathbf{a}(\mathbf{S};\boldsymbol{\theta})$, where $\boldsymbol{\theta}\in\boldsymbol{\Theta}$ and $\boldsymbol{\Theta}$ denotes a finite-dimensional set of parameters, and maximization is performed iteratively over the set $\boldsymbol{\theta}$. This leads to the parameterized multi-objective CPM problem
$\max_{\boldsymbol{\theta}\in\boldsymbol{\Theta}}{\frac{1}{T}\sum_{t=0}^{T-1}{f_{\text{scalar}}(\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t};\boldsymbol{\theta}))}},$ (5)
which can be solved through multi-objective RL algorithms. Although linear scalarization is computationally efficient and easy to implement, it assumes that the trade-offs between objectives can be accurately captured by the predefined weights in (4). In practice, selecting appropriate weight values can be challenging, especially when objectives exhibit nonlinear dependencies. More advanced multi-objective RL approaches attempt to overcome this limitation by dynamically adjusting the weighting scheme or learning a set of Pareto-optimal policies.
V Constrained Reinforcement Learning for CPM
As an alternative to multi-objective RL, the QoS requirements can be explicitly incorporated into a constrained RL problem. To formulate the CPM problem within constrained RL, we designate the first objective function in (3), $f_{0}(\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t}))$, as the primary objective to maximize, while treating the remaining objectives as QoS constraints that must be satisfied. The goal of the CPM problem is to determine the optimal CPM decision $\mathbf{a}(\mathbf{S}_{t})$ vector for any given network state $\mathbf{S}_{t}\in\mathpzc{S}$, such that the primary objective is optimized while ensuring compliance with QoS constraints. Accordingly, the generic parameterized constrained CPM problem is defined as
$\max_{\boldsymbol{\theta}\in\boldsymbol{\Theta}}\quad$ $\frac{1}{T}\sum_{t=0}^{T-1}f_{0}(\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t};\boldsymbol{\theta})),$ (6a)
s.t. $\frac{1}{T}\sum_{t=0}^{T-1}f_{i}(\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t};\boldsymbol{\theta}))\geq\text{c}_{i},\quad i=1,\dots,m-1,$ (6b)
where the constant $\text{c}_{i}\in\mathds{R}$ represents the threshold value for the $i$-th objective function. In this paper, we develop a learning algorithm to solve (6) for any given coexistence environment state $\mathbf{S}_{t}\in\mathpzc{S}$.
V-A Primal-Dual Approach for Constrained CPM
A customary approach to solve (6) is to consider a penalized version in the Lagrangian dual domain. Formally, we introduce dual variables $\boldsymbol{\lambda}\in\mathds{R}_{+}^{c}$ associated with the constraints in (6b) and define the Lagrangian
[…]
The Lagrangian in (7) should be maximized over $\boldsymbol{\theta}$, while subsequently minimizing over the dual variables $\boldsymbol{\lambda}$, i.e.,
[…]
$\boldsymbol{\theta}_{k}=\arg\max_{\boldsymbol{\theta}\in\boldsymbol{\Theta}}\left[\frac{1}{T_{0}}\sum_{t=kT_{0}}^{(k+1)T_{0}-1}f_{0}(\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t};\boldsymbol{\theta}))+\sum_{i=1}^{m-1}\lambda_{i,k}\left(\frac{1}{T_{0}}\sum_{t=kT_{0}}^{(k+1)T_{0}-1}\left(f_{i}(\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t};\boldsymbol{\theta}))-\text{c}_{i}\right)\right)\right],$ (9)
$\lambda_{i,k+1}=\left[\lambda_{i,k}-\frac{\eta_{\lambda_{i}}}{T_{0}}\sum_{t=kT_{0}}^{(k+1)T_{0}-1}\left(f_{i}(\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t};\boldsymbol{\theta}_{k}))-\text{c}_{i}\right)\right]^{+},\quad i=1,\dots,m-1.$ (10)
The advantage of replacing the objective in (6) with the Lagrangian in (7) is that the latter can be optimized using any parameterized learning framework, such as standard RL algorithms. One limitation of (8) is the ambiguity in determining suitable values for the dual variables. The optimal choice for $\boldsymbol{\lambda}$ depends on the transition probability $p(\mathbf{S}_{t+1}|\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t};\boldsymbol{\theta}))$, which is typically unknown. This challenge can be circumvented by dynamically adjusting the dual variables $\boldsymbol{\lambda}$. To accomplish this, we define an iteration index $k\in\left\{0,1,\dots,\left\lfloor T/T_{0}\right\rfloor-1\right\}$, where $T_{0}$ is the duration of one epoch, which is the number of time steps between successive updates of the model parameters. Also, we introduce the learning rates $\eta_{\lambda_{i}}\in\mathds{R}_{+}$ corresponding to each $\lambda_{i}$. The model parameters $\boldsymbol{\theta}_{k}$ and dual variables $\lambda_{i,k}$ are updated iteratively according to equations (9) and (10).
During the $k$-th epoch, the agent collects samples over a window of length $T_{0}$ and uses them to compute the constraint violation. The primal update in (9) seeks to maximize the Lagrangian, which includes the main objective $f_{0}$ and a penalty term weighted by the current dual variable $\lambda_{i,k}$ and the degree of constraint violations. The dual update in (10) adjusts each $\lambda_{i,k}$ based on the extent to which its corresponding constraint $\text{c}_{i}$ is violated. If the objective $f_{i}$ exceeds $\text{c}_{i}$, the dual variable increases proportionally to the violation, encouraging the agent to reduce the violation in future updates. The $[\cdot]^{+}$ operator ensures that the dual variables remain non-negative. The update rate $\eta_{\lambda_{i}}$ can be tuned individually for each constraint to control how aggressively violations are penalized.
It has been shown that the resulting algorithm is both feasible and near-optimal when run for a sufficiently large number of time steps. This is particularly noteworthy because, in standard primal-dual methods, the learned CPM policies are not guaranteed to yield a feasible set of decisions. In contrast, the CPM algorithm defined by equations (9) and (10) ensures feasibility and near-optimality. Moreover, this guarantee holds even when the underlying optimization problem is non-convex [8].
One of the important limitations of the primal-dual approach is that it only guarantees convergence to a feasible and near-optimal solution in the long run, i.e., as the total operation time $T$ becomes very large, theoretically approaching infinity. In practice, this means we cannot simply stop the algorithm after a finite number of iterations and claim that the solution it has reached is close to optimal. The performance and feasibility guarantees only hold in the limit, which may not be practical for systems that need to make decisions in real-time or within limited time frames.
Another challenge with this method is that optimizing the Lagrangian function at each iteration requires information about future network states. For example, when the algorithm reaches the beginning of the $k$-th epoch (at time step $t=kT_{0}$), the optimization process needs access to all network states from $t=kT_{0}$ to $t=(k+1)T_{0}-1$. However, in an online or real-time setting, this future information is not available, making it impossible to perform the optimization exactly as required. While this may be manageable in offline training where future states can be simulated or assumed, it poses a serious limitation during actual deployment.
V-B State-Augmented Constrained Reinforcement Learning
To address the aforementioned challenges, the state-augmentation approach provides a more practical and scalable alternative [8, 10, 9]. The key idea of state-augmentation is to treat constraint satisfaction as a dynamic component of the agent’s environment, which evolves in response to constraint violations through dual variable dynamics. By augmenting the state with dual variables, the agent learns policies that are constraint-aware and adaptable, leading to feasible, near-optimal solutions that traditional methods cannot guarantee.
State-augmentation approach reformulates the constrained RL problem by incorporating the dual variables directly into the agent’s state space. This simple yet powerful idea transforms the constrained optimization problem into a state augmented MDP, where the augmented state includes both the coexistence network’s state and the current values of the dual variables. By doing so, the agent is able to learn a single, adaptive policy that responds dynamically to changes in the dual variables, eliminating the need to optimize over multiple parameter sets or forecast future network states. Moreover, as discussed earlier, when dual variables evolve according to gradient-based updates, the learned policy remains feasible and near-optimal under general non-convex settings.
The state-augmentation approach offers several advantages over the classical Lagrangian-based method, where constraints are typically integrated into the reward function using fixed or static Lagrange multipliers. By embedding the dual variables directly into the agent’s state space and updating them dynamically, the agent is able to adaptively learn how to satisfy constraints over time. This dynamic adaptation is particularly beneficial in non-stationary or evolving environments, where fixed multipliers may fail to capture changing system dynamics. The advantage of state-augmentation approach becomes particularly significant in the context of CPM, where the environment is inherently dynamic and characterized by complex interactions between multiple heterogeneous transmitters where QoS requirements can vary rapidly. By continuously aligning coexistence parameter adjustments with the real-time status of QoS constraints, state augmentation helps ensure that critical requirements—such as bounded medium access delay for high-priority traffic—are upheld even in highly variable environments. This makes it particularly well-suited for managing the multi-objective and time-sensitive nature of CPM.
VI Proposed Algorithm: QaSAL-CPM
In light of the state-augmentation approach, we propose the QoS-aware State-Augmented Learnable algorithm for the CPM problem (QaSAL-CPM). Let us consider state $\mathbf{S}_{t}$ at time step $t$ of the $k$-th epoch. Augmentation of the dual variables $\boldsymbol{\lambda}_{k}$ into the state space results in a augmented state $\tilde{\mathbf{S}}_{t}=(\mathbf{S}_{t},\boldsymbol{\lambda}_{k}$). To be distinguished from previously introduced parameterization in Section V, we introduce an alternative parameterization for the CPM policy, in which the CPM decisions are represented via the parameterization $\mathbf{a}(\mathbf{\tilde{S}};\boldsymbol{\tilde{\theta}})$, where $\boldsymbol{\tilde{\theta}}\in\boldsymbol{\tilde{\Theta}}$ denotes the set of parameters of the state-augmented CPM policy. Then, we define the augmented version of the Lagrangian in (7) as
[…]
Figure 3: Flowchart of QaSAL-CPM: a) Training phase b) Execution phase.
Considering a probability distribution $p_{\boldsymbol{\lambda}}$ for the dual variables, we define the optimal state-augmented CPM policy which maximizes the expected augmented Lagrangian over this distribution as
[…]
Using the augmented policy parameterized by (12), we can obtain the CPM decisions $\mathbf{a}(\mathbf{\tilde{S}};\boldsymbol{\tilde{\theta}})$ which maximize the Lagrangian corresponding to the dual variables $\boldsymbol{\lambda}$ at each iteration $k$. The dual variable update equation in (10) should also be replaced with its augmented version:
$\lambda_{i,k+1}=\left[\lambda_{i,k}-\frac{\eta_{\lambda_{i}}}{T_{0}}\sum_{t=kT_{0}}^{(k+1)T_{0}-1}\left(f_{i}(\mathbf{\tilde{S}}_{t},\mathbf{a}(\mathbf{\tilde{S}}_{t};\boldsymbol{\tilde{\theta}}_{k}^{*}))-\text{c}_{i}\right)\right]^{+},$ (13)
where $i=1,\dots,m-1$.
To ensure the effectiveness of this approach, the chosen parameterization must possess sufficient expressive power. Specifically, the CPM decisions generated by (12) should be able to closely approximate those that would be obtained by solving the original constrained optimization problem iteratively for each $\lambda$, where the dependence of $\boldsymbol{\theta}_{k}$ and $\boldsymbol{\lambda}_{k}$ is made explicit, i.e., (9) and (10). This requirement places the chosen parameterization within a class known as near-universal parameterizations, which are capable of approximating the optimal mapping from dual variables to policy decisions with high fidelity. Examples of such expressive function approximators include neural networks and radial basis functions. These parameterizations are particularly effective in the constrained RL setting where the policy must adapt dynamically to changes in dual variables. To optimize such policies, standard RL algorithms, such as value-based methods or policy-gradient-based approaches, can be employed.
While these methods do not guarantee convergence to a globally optimal solution, they have demonstrated strong empirical performance and often converge to policies with minimal sub-optimality, especially when combined with expressive function approximators [10]. An important point is that the optimal model parameters $\boldsymbol{\tilde{\theta}}^{*}$ are directly utilized in (13), which mitigates the challenge posed in (9), where computing and storing the optimal set of parameters for every possible configuration of dual variables would be required. This was, in fact, the primary motivation for adopting the state-augmented parameterization approach.
VI-A Deep Reinforcement Learning for QaSAL-CPM
A flowchart depicting the QaSAL-CPM framework is given in Fig. 3. We apply a Double Deep Q-Network (DDQN), which decouples action selection and action evaluation during target computation by using the main Q-network to select the action and the target network to evaluate its value. This modification results in more accurate value estimates and improved learning stability, particularly in complex, dynamic environments. The training phase begins by initializing the model and target network parameters, as well as an empty experience replay memory. For each training episode, the environment is initialized, and a set of dual variables representing QoS constraints is sampled. These dual variables are then concatenated with the environment state to form an augmented state. The agent uses this augmented state to select a coexistence action that balances primary performance objectives with constraint satisfaction. After executing the action, the environment provides a reward that incorporates both the primary objective and penalties for constraint violations, and the agent transitions to a new augmented state. Each transition is stored in the replay memory for later training.
During training, the agent samples mini-batches from the replay memory and applies the DDQN mechanism to compute stable target Q-values. Specifically, the action is selected using the current Q-network, but its value is estimated using the target network. The algorithm then computes the augmented Lagrangian loss, which reflects both reward maximization and constraint handling, and updates the model parameters through gradient descent. The target network is periodically synchronized with the main Q-network to maintain training stability. This process is repeated over multiple episodes and time steps until convergence, ultimately yielding a trained model capable of making real-time CPM decisions that adhere to both performance and QoS constraints.
At each time step of the execution phase, the agent observes the current network state and augments it with the current dual variables to form the input for decision-making. Using the optimized model parameters from training, the agent makes a CPM decision that aims to optimize performance while meeting QoS requirements. Every $T_{0}$ steps, the algorithm updates the dual variables based on the average constraint violations observed over the last epoch according to (13). This update mechanism allows the agent to dynamically adjust its behavior in response to evolving constraint pressures, maintaining a balance between primary objectives and QoS guarantees throughout deployment. The training and execution methodologies for QaSAL-CPM utilizing the DDQN with experience replay are summarized in Algorithms 1 and 2, respectively.
Algorithm 1 Training phase of QaSAL-CPM with DDQN.
Input: Number of training episodes $N$, number of time steps $T$, QoS constraints $\mathbf{c}$, target network update frequency $\tau$, batch size $B$, replay memory capacity $M$, primal learning rate $\eta_{\boldsymbol{\tilde{\theta}}}\in(0,1]$, discount factor $\gamma\in(0,1]$.
Output: Optimal model parameters $\boldsymbol{\tilde{\theta}}^{*}$.
1:  Initialize:  Model parameters $\boldsymbol{\tilde{\theta}}_{0}$.  Target parameters $\boldsymbol{\tilde{\theta}}^{\text{target}}\leftarrow\boldsymbol{\tilde{\theta}}_{0}$.  Experience replay memory $\mathpzc{D}\leftarrow\varnothing$.
2:  for $n=0,1,\dots,N-1$ do
3:   Observe the initial network state $\mathbf{S}_{0}$.
4:   Randomly sample dual variables $\boldsymbol{\lambda}_{n}=\left\{\lambda_{i,n}\sim p_{\lambda}\right\}_{i=1}^{m-1}$.
5:   for $t=0,1,\dots,T-1$ do
a:
Augment network state $\mathbf{\tilde{S}}_{t}=(\mathbf{S}_{t},\boldsymbol{\lambda}_{n})$.
b:
Generate CPM decision $\mathbf{a}_{t}=\mathbf{a}(\mathbf{\tilde{S}}_{t};\boldsymbol{\tilde{\theta}}_{n})$.
c:
Calculate main reward $r_{t}=f_{0}(\mathbf{\tilde{S}}_{t},\mathbf{a}_{t})$.
d:
Calculate violations $v_{t}=\sum_{i=1}^{m-1}{\lambda_{i,n}\left(f_{i}(\mathbf{\tilde{S}}_{t},\mathbf{a}_{t})-\text{c}_{i}\right)}$.
e:
Observe and augment next state $\mathbf{\tilde{S}_{t+1}}=\left(\mathbf{S}_{t+1},\boldsymbol{\lambda}_{n}\right)$.
f:
Store transition $\left(\mathbf{\tilde{S}}_{t},\mathbf{a}_{t},r_{t},v_{t},\mathbf{\tilde{S}}_{t+1}\right)$ in $\mathpzc{D}$.
g:
Sample minibatch $\left\{\left(\mathbf{\tilde{S}}_{j},\mathbf{a}_{j},r_{j},v_{j},\mathbf{\tilde{S}}_{j+1}\right)\right\}_{j=1}^{B}$ from $\mathpzc{D}$.
h:
Calculate $\left\{y_{j}=r_{j}+\gamma\max_{\mathbf{a}^{\prime}}{Q(\mathbf{\tilde{S}}_{j+1},\mathbf{a^{\prime}}\mid\boldsymbol{\tilde{\theta}}_{n}^{\text{target}})}\right\}_{j=1}^{B}$.
i:
Compute the augmented Lagrangian loss:    $L(\boldsymbol{\tilde{\theta}}_{n})=\frac{1}{B}\sum_{j=1}^{B}{\left(y_{j}+v_{j}-Q(\mathbf{\tilde{S}}_{j},\mathbf{a}_{j}\mid\boldsymbol{\tilde{\theta}}_{n})\right)^{2}}$.
j:
Update model parameters via gradient descent:    $\boldsymbol{\tilde{\theta}}_{n+1}\leftarrow\boldsymbol{\tilde{\theta}}_{n}-\eta_{\boldsymbol{\tilde{\theta}}}\nabla_{\boldsymbol{\tilde{\theta}}_{n}}{L(\boldsymbol{\tilde{\theta}}_{n})}$.
k:
Update $\boldsymbol{\tilde{\theta}}^{\text{target}}\leftarrow\boldsymbol{\tilde{\theta}}_{n+1}$ every $\tau$ steps.
6:   end for
7:  end for
8:  $\boldsymbol{\tilde{\theta}}^{*}\leftarrow\boldsymbol{\tilde{\theta}}_{N}$
Algorithm 2 Execution phase of QaSAL-CPM.
Input: Optimal model parameters $\boldsymbol{\tilde{\theta}}^{*}$, dual variables update rates $\boldsymbol{\eta}_{\lambda}$, QoS constraints $\mathbf{c}$, epoch duration $T_{0}$.
Output: Sequence of CPM decisions $\{\mathbf{a}_{t};t=0,1,...\}$.
1:  Initialize: $\boldsymbol{\lambda}_{0}\leftarrow\boldsymbol{0}$, $k\leftarrow 0$.
2:  for $t=0,1,...$ do
a:
Observe and augment the state $\mathbf{\tilde{S}_{t}}=\left(\mathbf{S}_{t},\boldsymbol{\lambda}_{k}\right)$.
b:
Generate CPM decision $\mathbf{a}_{t}=\mathbf{a}(\mathbf{\tilde{S}}_{t};\boldsymbol{\tilde{\theta}}^{*})$.
c:
if $t+1\mod T_{0}=0$ then    Update the dual variables $\left\{\lambda_{i}\right\}_{i=1}^{m-1}$ as:     $\lambda_{i,k+1}=\left[\lambda_{i,k}-\frac{\eta_{\lambda_{i}}}{T_{0}}\sum_{t=kT_{0}}^{(k+1)T_{0}-1}\left(f_{i}(\mathbf{\tilde{S}}_{t},\mathbf{a}_{t})-\text{c}_{i}\right)\right]^{+}$.     $k\leftarrow k+1$.
d:
end if
3:  end for
VII Performance Evaluation and Simulation Results
In this section, we evaluate the proposed CPM algorithms in a realistic coexistence scenario involving 5G NR-U and Wi-Fi systems sharing an unlicensed spectrum. We implemented a Python simulation environment based on SimPy that models the MAC-layer behavior of 5G NR-U and Wi-Fi under saturated traffic conditions [^1] https://github.com/mfasihi/QaSAL_Coexistence_MAC. The simulated wireless environment is inherently dynamic and competitive due to the heterogeneous nature of the technologies involved. While 5G NR-U relies on scheduled transmissions with LBT, Wi-Fi uses a distributed contention-based access mechanism (i.e., CSMA/CA). These differences can lead to severe performance disparities and unpredictable behavior when the medium becomes congested. To investigate the behavior of the proposed methods, we first apply the multi-objective RL based algorithm from Section IV, which we call MORL-CPM, to a representative coexistence setting to demonstrate its limitations in handling competing objectives under hard QoS constraints. Subsequently, we apply the proposed QaSAL-CPM algorithm from Section VI across two distinct scenarios:
  • Scenario 1: A single gNB PC1 transmitter coexists with a fixed set of 25 PC3 transmitters from both NR-U and Wi-Fi networks.
  • Scenario 2: The gNB PC1 transmitter coexists with a symmetric and varying number of PC3 transmitters from both gNB and AP sides, ranging from 0 to 50.
VII-A Performance Metrics
We focus on two key performance metrics to evaluate the effectiveness of CPM: medium access delay and airtime fairness. Medium access delay is a critical factor for ensuring low-latency communication, particularly for high-priority traffic, while Jain’s fairness index provides a quantitative measure of how equitably resources are distributed among competing nodes. The two metrics are elaborated below.
VII-A1 Medium Access Delay
In coexistence scenarios where 5G NR-U and Wi-Fi share unlicensed spectrum, medium access delay plays a crucial role in ensuring Quality of Service (QoS), particularly for high-priority traffic such as URLLC, which demands stringent latency guarantees, often requiring end-to-end delays as low as $1$ ms with high reliability. Any excessive delay in accessing the channel or transmitting packets can lead to QoS degradation, packet drops, or violations of reliability constraints, making medium access delay a critical metric in CPM. By dynamically adjusting the contention parameters, the medium access delay can be controlled to ensure that high-priority URLLC traffic meets its latency requirements while maintaining fairness for other coexisting users.
In this paper, we define medium access delay as the time interval between the completion of a node’s successful transmission and the start of its next successful transmission. This delay includes backoff delay, which accounts for the time spent in the exponential backoff process before attempting transmission, and contention delay, which encompasses the total time a node waits due to channel occupancy by other devices. Additionally, when a collision occurs, the node must restart the contention process, further increasing the delay before the next successful transmission. Fig. 4 illustrates the calculation of medium access delay across the learning time steps. The cumulative medium access delay, $D_{t}$, is experienced by a node up to the end of time step $t$, accounting for the delays incurred in successful transmissions and collisions.
In an NR-U/Wi-Fi coexistence scenario the delay experienced by a high-priority PC1 transmitter is strongly influenced by the activity of other coexisting nodes, particularly lower-priority transmitters like those in PC3. These nodes often operate with fixed Maximum Channel Occupancy Time (MCOT), such as $8$ ms for PC3, which means that when a PC1 transmitter has to wait for the channel, its delay tends to increase in discrete steps. For instance, the delay could jump to 8 ms, 16 ms, or more, depending on how many consecutive MCOT periods it is blocked. As a result, the delay values do not vary smoothly but appear in distinct levels. This makes it difficult to access performance based on a single delay value, since it might be unusually high due to transient interference or contention.To address this, we use a smoothed medium access delay, $\overline{D}_{t}$, which averages the last few delays (e.g., the last five) observed for the transmitter. This helps reduce the sensitivity to transient spikes caused by contention or collision, and provides a more consistent reflection of ongoing latency performance. As a result, the learning algorithm can make more informed and stable decisions when optimizing coexistence under QoS constraints.
VII-A2 Airtime Fairness
While reducing medium access delay is critical for high-priority traffic, it must be balanced with fair resource allocation between PC1 and PC3 transmitters from both the NR-U and Wi-Fi networks. An overly aggressive optimization for PC1 traffic could starve lower-priority traffic, leading to unfair spectrum access and potential performance degradation. To prevent this, fairness should be incorporated as a key metric in CPM optimization. Jain’s Fairness Index (JFI) is a widely used metric to assess the fairness of resource allocation among competing entities. Considering the PC1 and PC3 transmitters as these entities, we define JFI as follows:
$\text{JFI}=\frac{\left(\text{Airtime}_{\text{ PC1}}+\text{Airtime}_{\text{ PC3}}\right)^{2}}{2\left(\text{Airtime}_{\text{ PC1}}^{2}+\text{Airtime}_{\text{ PC3}}^{2}\right)}.$ (14)